OpenStudy (anonymous):

Simplify x squared minus 12 times x plus 20 all over x minus 10. A:x − 2 B:x minus 2 all over x minus 10 C:x minus 2 all over 10 D:x minus 10 all over x minus 2

OpenStudy (anonymous):

a

OpenStudy (anonymous):

\[\frac{x^2 - 12x + 20}{x-10} = \frac{\cancel{(x-10)} \times (x-2)}{\cancel{(x-10)}} = (x - 2)\]